the width of a rectangle is 1/3 of its length. if the perimeter is 96 meters. find its length
Answer by user_dude2008(1862)

"the width of a rectangle is 1/3 of its length" --> W=(1/3)L


P=2L+2W

96=2L+2(1/3)L

96=2L+(2/3)L

96=(8/3)L

L=96(3/8)

L=36 <---- length